Circular No. 7102

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD P. Garnavich, R. Kirshner, and P. Challis, Harvard-Smithsonian Cen-

ter for Astrophysics, report on behalf of the Supernova Intensive Study team: “Images of SN 1987A were obtained with the Wide Field/Planetary Camera II (WFPC2) onboard the Hubble Space Telescope on Jan. 7–8. The data show that the ring ‘hotspot’ (cf. IAUC 6665, 6710) is continuing to brighten at wavelengths between 200 and 900 nm. The hotspot flux in the F656N filter (Hα + N II 654.8-nm) is now 6.4 × 10^{−15} erg cm^{−2} s^{−1}, and the F675W magnitude (similar to R band) is 20.65 ± 0.05. The F675W light curve over the past four years is well fitted by a line with a slope of −0.0024 mag/day, more than doubling in brightness each year. The hotspot is now easily detected in the F255W filter, with a flux of ∼ 3 × 10^{−15} erg cm^{−2} s^{−1}. The new F675W image clearly shows evidence of the reverse-shock that had been detected spectroscopically in Lyα and Hα (Michael et al. 1998, Ap.J. 509, L117). The limb-brightened shock appears as two arcs within the ring, east and west of the debris. Their separation is measured to be 1′′.20 ± 0′′.04 along the major axis of the ring. Assuming spherical expansion and a maximum observed velocity of the Lyα emission of −12000 ± 1400 km/s (Michael et al., with the error based on the slit width), the distance to SN 1987A is 50 ± 6 kpc. This is consistent with most distance estimates to the LMC. The east arc is clearly brighter than the western arc by a factor of ∼ 1.5, mirroring the asymmetry found in the radio by Gaensler et al. (1997, Ap.J. 479, 845) and calling into question the assumption of spherical expansion.”

3C 279 M. Villata, C. M. Raiteri, G. Sobrito, G. De Francesco, and L. Lanteri,

Osservatorio Astronomico di Torino; and R. C. Hartman, Goddard Space Flight Center, for the CGRO/EGRET Team, write: “3C 279 is optically bright and active; CCD data taken with the Torino 1.05-m telescope from Jan. 14 to 26 show fast variations in R in the range 13.78–14.50 (± 0.02; see

http://www.to.astro.it/Groups/Extragal/blazars.htm). Preliminary

results from EGRET observations show a flux (> 100 MeV) of ∼ 2 × 10^{−6} photon cm^{−2} s^{−1}, with some evidence for time variation over the first 4.5 days (Jan. 20.5–24.0 UT). From preliminary analysis of data from RXTE, A. Lawson and I. McHardy report that the x-ray flux is at an intermediate historical level, consistent with the present EGRET flux. The RXTE PCA countrate (3–10 keV) is very roughly 3 counts/s for 3 PCUs.”

1999 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green